OpenStudy (anonymous):

Solve the equation (eqn coming soon) . . . ANSWER = {12}

OpenStudy (anonymous):

\[\sqrt{3x-11}+\sqrt{3x}=\sqrt{12x-23}\]

OpenStudy (anonymous):

hello anyone there

OpenStudy (anonymous):

okay??

OpenStudy (anonymous):

thanks alot

OpenStudy (anonymous):

but u could have given hints

OpenStudy (anonymous):

so i could solve it my self

OpenStudy (anonymous):

sorry..will keep in mind next time..:)

OpenStudy (anonymous):

but thanks it was pretty easy, if you had seen mine register u would fall from your seat

OpenStudy (anonymous):

it was a mess!!

OpenStudy (anonymous):

it happens..mistakes are the most significant part of learning..:)

OpenStudy (anonymous):

but i do those mistakes alot some people here in openstudy know that ;)